# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ IV.1 Model systems
While discussing various applications of the developed stochastic calculus for MJP we will throughout elude to biophysically relevant model systems shown in Fig. 6. Thereby we aim mostly at illustrating how to apply the theoretical results rather than to discuss the respective biological implications.
Figure 6: Overview of models used to visualize the results: (a) Markov model for secondary active transport (SAT); the states $(1,3,5)$ correspond to the protein funnel being open toward the cell exterior while states $(2,4,6)$ describe the funnel being open toward the cell interior. The transition rate for the conversion of the opening from the interior to the exterior (and vice versa) is $\gamma$, which in the presence of ligands AA or $B$ is scaled by a dimensionless factor $x$. AA and $B$ leave the funnel with rate $l_{A/B}$ and enter it with rate ${\rm e}^{\mathrm{out}/\mathrm{in}}_{A/B}$ outside/inside the cell. Of particular interest in the (extended) SAT model is the impact of ${\rm e}^{\mathrm{out}}_{B}$, i.e., the concentration of the helper molecule. (b) Schematic of how the SAT model in (a) may be extended to a “train” of SAT models to capture the number of AA molecules, $n$, transported through the cell membrane. (c) Schematic of the Markov model for calmodulin folding. We focus on the inference of dissipation by observing a current with weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{j2}\delta_{i3}$ and/or density with weight $V_{i}=(\delta_{i1}+\delta_{i2})/2$. (d) Four-state ring model with equilibrium transition rates parameterized by free energies $E_{i}$ for state $i$, local diffusion coefficient $D$, temperature $T$, and mixing parameter $\lambda$ [see Eq. (63)], and upon constant driving, respectively. We are interested in the response to temperature perturbations and investigate the effect on a single-time observable $A_{t}=\delta_{x_{t}1}+\delta_{x_{t}2}$ and a current with transition weights $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$. (e) 4-state ring model with constant driving having rates 1 in the clockwise and $\epsilon$ in counterclockwise direction; the stationary distribution is independent of $\epsilon$. The model is used for a comparison of the quality of thermodynamic bounds in situations where only the mesostates A and B, corresponding to states $(1,2)$ and $(3,4)$, respectively, are observed without the ability to distinguish the respective transition pathways between them.
@@ IV.1.1 Secondary Active Transport
We first consider a Markov model of secondary active transport (SAT) of molecule $A$ through a protein funnel facilitated by molecule $B$ [145, 146, 147]. The model, shown in Fig. 6a, consists of six states; three [i.e., $(1,3,5)$] corresponding to the funnel being open toward the exterior of the cell and other three [i.e., $(2,4,6)$] open toward the interior of cell. In the absence of $A$ or $B$ bound, the funnel changes conformation with rate $\gamma$. Conversely, the crossing via $A$ or $B$-filled states occurs with $\gamma x$ for a dimensionless speedup factor $x$. AA and $B$ leave or enter the funnel inside and outside the cell with rates $l_{A/B}$ and ${\rm e}^{\mathrm{out/in}}_{A/B}$, respectively.
We further consider a periodic extension of coupled SAT models to describe the transport of molecules $A$ through the membrane (Fig. 6b). We leave the rates in the system unchanged, whereas the state space is extended to $\mathcal{N}\times\mathds{Z}$, with the additional degree of freedom accounting for the total number of $A$ molecules outside the cell.
@@ IV.1.2 Calmodulin Folding Dynamics
Being a $\mathrm{Ca}^{2+}$ receptor, calmodulin plays an important role in physiological signaling processes [148]. The Markov model for the calmodulin folding pathway [88] is commonly used as a paradigmatic example in, e.g., the study of first passage times [149] and stochastic thermodynamics [59, 150]. The respective states represent different conformations of the protein and the transitions are changes of conformation [88]. Here we adopt the transition rates from Ref. [149] summarized in Tab. 2 while the topology of the folding network is shown in Fig. 6c. The calmodulin example will be used to gauge the quality of thermodynamic inference via the derived thermodynamic inequalities.
@@ IV.1.3 Four-State Ring Toy Model
We use the free energies listed in Tab. 4.
@@ IV.2.1 Entropy and Pseudo-Entropy Production
$A_{t}^{\mathrm{c}}\equiv\int_{\tau=0}^{\tau=t}\frac{\bm{j}(\bm{x}_{\tau},\tau)}{P(\bm{x}_{\tau},\tau)}\left[2\bm{D}(\bm{x}_{\tau})\right]^{-1}\bm{\sigma}(\bm{x}_{\tau})\mathrm{d}\bm{W}_{\tau},$ (65)
@@
$A_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[z_{\tau}^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (67)
@@ IV.2.2 General TUR and Correlation TUR
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (75)
@@
$\tilde{\rho}_{t}$ $\equiv\int_{\tau=0}^{\tau=t}\tau\partial_{\tau}\bm{V}(\tau)^{T}\mathrm{d}\bm{\tau}(\tau).$ (79)
By shifting the current $\Delta J_{t}\to\Delta J_{t}-c(t)\Delta\rho_{t}$ with some $c:\mathds{R}\to\mathds{R}$ in Eq. (74) already yields the transient correlation TUR (CTUR)
@@ IV.2.3 Optimization of TUR and CTUR
There are two ways to optimize Eq. (80), either by saturating Eq. (74) or by choosing the optimal $c(t)$ for fixed $J_{t}$ and $\rho_{t}$. The former requires $\Delta J_{t}-c(t)\Delta\rho_{t}\propto A_{t}$. Since $c(t)\neq 0$, we see that the saturation is achieved when $V_{i}(\tau)=\hat{\bm{e}}_{i}^{T}\bm{\kappa}(\tau)\bm{L}\hat{\bm{e}}_{i}/c(t)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$ and $\bm{\kappa}=c^{\prime}\bm{Z}$ for any $c^{\prime}\in\mathds{R}$. Here, $\hat{\bm{e}}_{i}$ is the $i$th unit vector. However, these choices do not saturate Eq. (80) but only the estimate for the pseudo-entropy production [see first inequality in Eq. (78)]. In other words, the “saturated” CTUR quality factor reads
@@
$c^{\dagger}(t)$ $=\frac{a(t)}{b(t)}\,$,
where $a(t)\equiv t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle$ and $b(t)\equiv(t\partial_{t}-1)\langle{\rho}_{t}\rangle-\langle{\tilde{\rho}}_{t}\rangle$. All quantities entering Eq. (83) are accessible from $J_{t}$ and $\rho_{t}$, making the optimization practically feasible. Note that $c^{\dagger}(t)$ and $c^{*}(t)$ correspond to the minimum and maximum of $h(t)$, respectively, as
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{\dagger}(t)}$ $=\frac{4b(t)^{2}}{\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}\geq 0\,$, (84)
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{*}(t)}$ $=\frac{-4b(t)^{2}\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}{\left(\mathrm{var}\left({J}_{t}-c^{*}(t)\rho_{t}\right)\right)^{2}}\leq 0\,.$
Moreover, one can immediately see that using $c^{\dagger}(t)$ one recovers the second law of thermodynamics, as $t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c^{\dagger}(t)\left[(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right]=0$. The optimal quality factor therefore reads
@@
The saturated quality factor $\mathcal{Q}_{\mathrm{sat}}^{\mathrm{CTUR}}$ in Eq. (82) is shown in Fig. 7 together with the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ and the quality factors of the current TUR $\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$, density TUR $\mathcal{Q}^{\rho\mathrm{TUR}}_{12}$, and a CTUR $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}$ with $c(t)=10$ for the calmodulin system (see Fig. 6c). We consider the transition weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$ and state function $V_{i}=(\delta_{i1}+\delta_{i2})/2$. Only the saturated quality factor remains finite for $t\to\infty$. As expected, the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ gives the best inferred entropy production when $\bm{\kappa}(\tau)$ and $\bm{V}(\tau)$ do not saturate the TUR, whereas $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}\leq\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$ for $t\leq 3$. Hence, simply using an arbitrary $c(t)$ in the CTUR may in fact lead to a worse estimate for the entropy production compared to the current TUR. The non-trivial behavior of $c(t)$ is also highlighted in Fig. 8. The expressions of $c^{*}(t)$ and $c^{\dagger}(t)$ Eqs. (83) can be both positive and negative. Furthermore, they are in no particular relation to another, i.e., one cannot easily say if and when $c^{*}(t)<c^{\dagger}(t)$ and vice versa. A more detailed discussion is given in App. A.10.
@@
Figure 7: Saturation of CTUR in the transient calmodulin folding dynamics (see Fig. 6c). We present the various CTUR quality factors $\mathcal{Q}$ arising from various combinations of density, $\rho$, and current, $J_{t}$, as a function of time. The black line denotes the saturation of the TUR, i.e., when $\bm{\kappa}(\tau)=z_{\tau}$ and $c(t)\rho_{t}=J_{t}^{\mathrm{II}}$, corresponding to the estimated entropy production being the pseudo-entropy production $\Sigma_{\mathrm{est}}=2\langle A_{t}^{2}\rangle$. Above the black line is the region $\Sigma_{\mathrm{est}}>2\langle A_{t}^{2}\rangle$ which is not accessible using the (C)TUR. The remaining quality factors shown use the current ($\mathcal{Q}_{(23)}^{J\mathrm{TUR}}$), density ($\mathcal{Q}_{12}^{\rho\mathrm{TUR}}$), or a combination of these ($\mathcal{Q}_{c=10}^{\mathrm{CTUR}}$ and $\mathcal{Q}_{\mathrm{opt}}^{\mathrm{CTUR}}$) in orange, green, purple, and red, respectively. For the CTURs, the proportionality $c(t)=10$ (purple) and $c(t)=c^{*}(t)$ (red) are chosen, the latter optimizing the quality factor for given $J_{t}$ and $\rho_{t}$. The rates can be found in Tab. 2.
@@
Figure 8: Quality factor as a function of density weighting function $c(t)$ in the calmodulin system at a fixed $t=0.05$. The color correspond to different transition weights $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$ between two states, i.e., $\kappa_{ij}^{(12)}$ (blue), $\kappa_{ij}^{(56)}$ (orange), and $\kappa_{ij}^{(26)}$ (green). The state function entering the density is $V_{i}(\tau)=\delta_{i3}$ and the initial condition is $p_{i}(0)=\delta_{i6}$. The circles are the quality factors for most optimal $c(t)=c^{*}(t)$, while the squares denote the least optimal $c(t)=c^{\dagger}(t)$.

### v2

@@ IV.1 Model systems
While discussing various applications of the developed stochastic calculus for MJP we will throughout allude to biophysically relevant model systems shown in Fig. 6. Thereby we aim mostly at illustrating how to apply the theoretical results rather than to discuss the respective biological implications. The three models we introduce in the following sections are the secondary active transport across cell membranes (and an extension of this model), the calmodulin folding dynamics, and a four-state toy model in various settings.
Figure 6: Overview of models used to visualize the results: (a) Markov model for secondary active transport (SAT); the states $(1,3,5)$ correspond to the protein funnel being open toward the cell exterior while states $(2,4,6)$ describe the funnel being open toward the cell interior (see Tab. 2). The transition rate for the conversion of the opening from the interior to the exterior (and vice versa) is $\gamma$, which in the presence of ligands AA or $B$ is scaled by a dimensionless factor $x$. AA and $B$ leave the funnel with rate $l_{A/B}$ and enter it with rate ${\color[rgb]{0,0,0}e}^{\mathrm{out}/\mathrm{in}}_{A/B}$ outside/inside the cell. Of particular interest in the (extended) SAT model is the impact of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$, i.e., the concentration of the helper molecule. (b) Schematic of how the SAT model in (a) may be extended to a “train” of SAT models to capture the number of AA molecules, $n$, transported through the cell membrane. (c) Schematic of the Markov model for calmodulin folding. We focus on the inference of dissipation by observing a current with weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{j2}\delta_{i3}$ and/or density with weight $V_{i}=(\delta_{i1}+\delta_{i2})/2$. (d) Four-state ring model with equilibrium transition rates parameterized by free energies $E_{i}$ for state $i$, local diffusion coefficient $D$, temperature $T$, and mixing parameter $\lambda$ [see Eq. (71)], and upon constant driving, respectively. We are interested in the response to temperature perturbations and investigate the effect on a single-time observable ${\color[rgb]{0,0,0}A}_{t}=\delta_{x_{t}1}+\delta_{x_{t}2}$ and a current with transition weights $\kappa_{ij}=\delta_{i4}\delta_{j1}-\delta_{i1}\delta_{j4}+\delta_{i3}\delta_{j2}-\delta_{i2}\delta_{j3}$. (e) 4-state ring model with constant driving having rates 1 in the clockwise and $\epsilon$ in countercockwise direction; the stationary distribution is independent of $\epsilon$. The model is used for a comparison of the quality of thermodynamic bounds in situations where only the mesostates A and B, corresponding to states $(1,2)$ and $(3,4)$, respectively, are observed without the ability to distinguish the respective transition pathways between them.
@@ IV.1.1 Secondary Active Transport
Table 2: Overview of states in SAT model.
State Meaning
1 Funnel outside with $A$ bound
2 Funnel inside with $A$ bound
3 Funnel outside with nothing bound
4 Funnel inside with nothing bound
5 Funnel outside with $B$ bound
6 Funnel inside with $B$ bound
We first consider a Markov model of secondary active transport (SAT) of molecule $A$ through a protein funnel facilitated by molecule $B$ [14, 16, 15]. The model, shown in Fig. 6a and listed in Tab. 2, consists of six states; three [i.e., $(1,3,5)$] corresponding to the funnel being open toward the exterior of the cell and other three [i.e., $(2,4,6)$] open toward the interior of cell. Hence, the state space is $\mathcal{N}=\{1,2,3,4,5,6\}$. In the absence of $A$ or $B$ bound, the funnel changes conformation with rate $\gamma$. Conversely, the crossing via $A$ or $B$-filled states occurs with $\gamma x$ for a dimensionless speedup factor $x$. AA and $B$ leave or enter the funnel inside and outside the cell with rates $l_{A/B}$ and ${\rm e}^{\mathrm{out/in}}_{A/B}$, respectively. The adopted parametrization is listed in Tab. 4 in Appendix A.11.
We further consider a periodic extension of coupled SAT models to describe the total transport of molecules $A$ through the membrane (Fig. 6b). We leave the rates in the system unchanged, whereas the state space is extended to $\mathcal{N}\times{\color[rgb]{0,0,0}\mathbb{Z}}$, with the additional degree of freedom accounting for the total number of $A$ molecules outside the cell.
The transport can be split into two categories: symporters and antiporters. These refer to to molecule $A$ being transported parallel or anti-parallel to molecule $B$, respectively. We are interested in the fraction of the total steady-state entropy production that we can estimate via the thermodynamic bounds for various values of transition rates ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$. These transition rates are assumed proportional to the concentration of $B$ outside the cell [14], allowing us to “control” the value of ${\color[rgb]{0,0,0}e}^{\mathrm{out}}_{B}$, leaving it as free parameter we can tune.
@@ IV.1.2 Calmodulin Folding Dynamics
Being a $\mathrm{Ca}^{2+}$ receptor, calmodulin plays an important role in physiological signaling processes [155]. The Markov model for the calmodulin folding pathway [156] is commonly used as a paradigmatic example in, e.g., the study of first passage times [12] and stochastic thermodynamics [90, 34]. The respective states represent different conformations of the protein and the transitions are changes of conformation [156]. Here we adopt the transition rates from Ref. [12] summarized in Tab. 3 in Appendix A.11 while the topology of the folding network is shown in Fig. 6c. The calmodulin example will be used to gauge the quality of thermodynamic inference via the derived thermodynamic inequalities.
@@ IV.1.3 Four-State Ring Toy Model
We use the free energies listed in Tab. 5 in Appendix A.11.
@@ IV.2.1 Entropy and Pseudo-Entropy Production
$A_{t}^{\mathrm{c}}\equiv\int_{\tau=0}^{\tau=t}\frac{\bm{j}(\bm{x}_{\tau},\tau){\color[rgb]{0,0,0}{}^{T}}}{P(\bm{x}_{\tau},\tau)}\left[2\bm{D}(\bm{x}_{\tau})\right]^{-1}\bm{\sigma}(\bm{x}_{\tau})\mathrm{d}\bm{W}_{\tau},$ (73)
@@
$A_{t}$ $=\int_{\tau=0}^{\tau=t}\mathrm{Tr}[{\color[rgb]{0,0,0}\bm{Z}(\tau)}^{T}\mathrm{d}\bm{\varepsilon}(\tau)],$ (75)
@@ IV.2.2 General TUR and Correlation TUR
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$ (83)
@@
[…]
By shifting the current $\Delta J_{t}\to\Delta J_{t}-c(t)\Delta\rho_{t}$ with some $c:{\color[rgb]{0,0,0}\mathbb{R}}\to{\color[rgb]{0,0,0}\mathbb{R}}$ in Eq. (82) already yields the transient correlation TUR (CTUR)
@@ IV.2.3 Optimization of TUR and CTUR
There are two ways to optimize Eq. (88), either by saturating Eq. (82) or by choosing the optimal $c(t)$ for fixed $J_{t}$ and $\rho_{t}$. The former requires $\Delta J_{t}-c(t)\Delta\rho_{t}\propto A_{t}$. Since $c(t)\neq 0$, we see that the saturation is achieved when $V_{i}(\tau)=\hat{\bm{e}}_{i}^{T}\bm{\kappa}(\tau)\bm{L}\hat{\bm{e}}_{i}/c(t)=\sum_{j}\kappa_{ij}(\tau)r_{ij}/c(t)$ and $\bm{\kappa}=c^{\prime}\bm{Z}$ for any $c^{\prime}\in{\color[rgb]{0,0,0}\mathbb{R}}$. Here, $\hat{\bm{e}}_{i}$ is the $i$th unit vector. However, these choices do not saturate Eq. (88) but only the estimate for the pseudo-entropy production [see first inequality in Eq. (86)]. In other words, the “saturated” CTUR quality factor reads
@@
$c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ $=\frac{a(t)}{b(t)}\,$,
$c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ $=\frac{a(t)\mathrm{cov}({\rho}_{t},{J}_{t})-b(t)\mathrm{var}({J}_{t})}{a(t)\mathrm{var}({\rho}_{t})-b(t)\mathrm{cov}({\rho}_{t},{J}_{t})}\,$, (91)
where $a(t)\equiv t\partial_{t}\langle{J}_{t}\rangle-\langle{\tilde{J}}_{t}\rangle$ and $b(t)\equiv(t\partial_{t}-1)\langle{\rho}_{t}\rangle-\langle{\tilde{\rho}}_{t}\rangle$. All quantities entering Eq. (91) are accessible from $J_{t}$ and $\rho_{t}$, making the optimization practically feasible. Note that $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ correspond to the minimum and maximum of $h(t)$, respectively, as
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)}$ $=\frac{4b(t)^{2}}{\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}\geq 0\,$, (92)
$\frac{\partial^{2}}{\partial c^{2}}h(c,t)|_{c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)}$ $=\frac{-4b(t)^{2}\mathrm{var}\left({J}_{t}-\frac{a(t)}{b(t)}\rho_{t}\right)}{\left(\mathrm{var}\left({J}_{t}-c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)\rho_{t}\right)\right)^{2}}\leq 0\,.$
Moreover, one can immediately see that using $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ one recovers the second law of thermodynamics, as $t\partial_{t}\langle J_{t}\rangle-\langle\tilde{J}_{t}\rangle-c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)\left[(t\partial_{t}-1)\langle\rho_{t}\rangle-\langle\tilde{\rho}_{t}\rangle\right]=0$. The optimal quality factor therefore reads
@@
The saturated quality factor $\mathcal{Q}_{\mathrm{sat}}^{\mathrm{CTUR}}$ in Eq. (90) is shown in Fig. 7 together with the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ and the quality factors of the current TUR $\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$, density TUR $\mathcal{Q}^{\rho\mathrm{TUR}}_{12}$, and a CTUR $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}$ with $c(t)=10$ for the calmodulin system (see Fig. 6c). We consider the transition weights $\kappa_{ij}^{(23)}=\delta_{i2}\delta_{j3}-\delta_{i3}\delta_{j2}$ and state function $V_{i}=(\delta_{i1}+\delta_{i2})/2$. Only the saturated quality factor remains finite for $t\to\infty$. As expected, the optimal $\mathcal{Q}^{\mathrm{CTUR}}_{\mathrm{opt}}$ gives the best inferred entropy production when $\bm{\kappa}(\tau)$ and ${V}_{\tau}$ do not saturate the TUR, whereas $\mathcal{Q}^{\mathrm{CTUR}}_{c=10}\leq\mathcal{Q}^{J\mathrm{TUR}}_{(23)}$ for $t\leq 3$. Hence, simply using an arbitrary $c(t)$ in the CTUR may in fact lead to a worse estimate for the entropy production compared to the current TUR. The non-trivial behavior of $c(t)$ is also highlighted in Fig. 8. The expressions of $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ Eqs. (91) can be both positive and negative. Furthermore, they are in no particular relation to another, i.e., one cannot easily say if and when $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)<c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ and vice versa. A more detailed discussion is given in App. A.10.
@@
Figure 7: Saturation of CTUR in the transient calmodulin folding dynamics (see Fig. 6c). We present the various CTUR quality factors $\mathcal{Q}$ arising from various combinations of density, $\rho_{{\color[rgb]{0,0,0}\tau}}$, and current, $J_{t}$, as a function of time. The black line denotes the saturation of the TUR, i.e., when $\bm{\kappa}(\tau)={\color[rgb]{0,0,0}\bm{Z}(\tau)}$ and $c(t)\rho_{t}=J_{t}^{\mathrm{II}}$, corresponding to the estimated entropy production being the pseudo-entropy production $\Sigma_{\mathrm{est}}=2\langle A_{t}^{2}\rangle$. Above the black line is the region $\Sigma_{\mathrm{est}}>2\langle A_{t}^{2}\rangle$ which is not accessible using the (C)TUR. The remaining quality factors shown use the current ($\mathcal{Q}_{(23)}^{J\mathrm{TUR}}$), density ($\mathcal{Q}_{12}^{\rho\mathrm{TUR}}$), or a combination of these ($\mathcal{Q}_{c=10}^{\mathrm{CTUR}}$ and $\mathcal{Q}_{\mathrm{opt}}^{\mathrm{CTUR}}$) in orange, green, purple, and red, respectively. For the CTURs, the proportionality $c(t)=10$ (purple) and $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ (red) are chosen, the latter optimizing the quality factor for given $J_{t}$ and $\rho_{t}$. The rates can be found in Tab. 3.
@@
Figure 8: Quality factor as a function of density weighting function $c(t)$ in the calmodulin system at a fixed $t=0.05$. The color correspond to different transition weights $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$ between two states, i.e., $\kappa_{ij}^{(12)}$ (blue), $\kappa_{ij}^{(56)}$ (orange), and $\kappa_{ij}^{(26)}$ (green). The state function entering the density is $V_{i}(\tau)=\delta_{i3}$ and the initial condition is $p_{i}(0)=\delta_{i6}$. The circles are the quality factors for most optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$, while the squares denote the least optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$.